OpenStudy (anonymous):

alg 2 help? functions?

OpenStudy (anonymous):

this is what i got. The function f(x) is a difference of squares. ---- y = x^2 -9 The function g(x) is a sum of squares. ---- y = x^2 9 The function h(x) is a perfect square trinomial. ---- y = x^2 6x 9 The function j(x) can only have a GCF factored out of it. --- y = x^2 2x 1st y = (x 3)(x-3) 2nd is prime for real numbers 3rd y = (x 3)^2 4th y = x(x 2)

OpenStudy (anonymous):

my teacher said "actually show how you would factor each if possible. Also on the last two you are missing + and - symbols, please add so we can grade accurately"

OpenStudy (anonymous):